1.There is a fruit seller who has 30 Kgs of apples to be sold and he wants to fix a price so that all the apples are sold. There are three customers in the market and their individual demand functions are given below:

D1=25-.05P
D2=20-.025P
D3=15-.075P
Where D is the demand and P is the price

Determine
Market demand equation for the fruit seller
Price at which he can sell all the apples
Individual demands of each of the three customers

2. A) Determine the market equilibrium price if the demand and supply function is given as:
D = 12p + 8
S = 14p – 4

Where D= demand
S=supply
p= price

B) Determine the equilibrium quantity if price is the same as above
D = 4p – 4q
S = 8q – 4p
Where D= demand S=supply p=price q=quantity
1
Expert's answer
2015-09-04T00:00:42-0400
1) If Qs = 30 Kgs of apples to be sold and
there are three customers in the market and their individual demand functions are given below:
D1=25-.05P
D2=20-.025P
D3=15-.075P
Then the market demand equation for the fruit seller is: D = D1 + D2 + D3 = 60 ¬ 0.15P
D = S, so:
60 - 0.15P = 30,
0.15P = 30,
P = $200 is the price at which he can sell all the apples.
The individual demands will be:
D1 = 25 - .05*200 = 15 units
D2 = 20 - .025*200 = 15 units
D3 = 15 - .075*200 = 0 units
2) a) D = 12p + 8
S = 14p – 4
Equilibrium price is in the point, where Qd = Qs, so:
12p + 8 = 14p - 4
p = $6
b) The equilibrium quantity is in the point, where Qd = Qs, so if p = 6, then:
D = 4p – 4q
S = 8q – 4p
24 - 4q = 8q - 24
q = 4 units.